2 mystery numbers=138. One of the numbers is 32 more than the other. What are the 2 numbers
Answer by dfrazzetto(283)


x+y = 138
x = y + 32
y + 32 + y = 138
2y + 32 = 138
2y = 106
y= 53
x= 85